# A Machine Learning and Finite Element Framework for Inverse Elliptic PDEs via Dirichlet-to-Neumann Mapping

### v1

@@ 3.4 Loss Functions
The loss function based on the Problem is defined as
@@
where $\mathcal{D}=\{(g_{i},h_{g_{i}})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\text{ supp }g_{i}\subset\omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\omega}\}$, $\tilde{\mathcal{D}}=\{(g_{i},\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}|_{\omega})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\text{ supp }g_{i}\subset\omega\}$, and $\|\cdot\|_{L^{2}(\omega)}$ denotes the $L^{2}$ norm on $\omega$. The loss function $\mathcal{L}$ is minimized by adjusting the model parameters $\boldsymbol{\theta}$. When the loss function reaches its minimum, we have $\mathcal{D}=\tilde{\mathcal{D}}$.
@@
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}}=\frac{\partial\mathcal{L}}{\partial k_{\text{fem}}}\cdot\frac{\partial\tilde{k}}{\partial\boldsymbol{\theta}}.$ (4)
@@
4 Numerical Experiments
In this section, we present numerical experiments to validate and demonstrate the effectiveness of the proposed framework introduced in Section 3. The following setup is common across all examples. The entire dataset consists of 4,096 samples: 2,048 for training, 1,024 for validation, and 1,024 for testing. The neural network is trained for 100 epochs with a batch size of 1. The learning rate is set to 0.001 for the first 50 epochs and reduced to 0.0001 for the remaining 50 epochs to facilitate stable convergence and prevent overshooting during the later stages of training. Training is performed using PyTorch [19] and the ADAM optimizer [20], a variant of stochastic gradient descent. All experiments are conducted on an NVIDIA GeForce RTX 3090 GPU (24GB VRAM), an Intel Core i9-12900K CPU, and 62GB RAM. The system runs with CUDA 11.8 and NVIDIA driver version 522.06.
In the next subsections, we evaluate the performance of the proposed method in several cases:
4.1 Example 1: Constant Coefficient
In this example, we consider a constant coefficient $k$ to provide a baseline for evaluating model performance under simple conditions. Specifically, we set $k=1$. The computational domain is a unit square $\Omega=(0,1)\times(0,1)$ represented by a mesh with 66,049 points. The Dirichlet boundary condition is given by $g=\cos(a\pi x)\cos(b\pi y)|_{\partial\Omega}$, where, $a$ and $b$ are independently sampled from a uniform distribution over $[0,2]$. The source function is defined as $f(x,y)=(a^{2}+b^{2})\pi^{2}\cos(a\pi x)\cos(b\pi y)$.
To evaluate how effectively our method reconstructs the coefficient $\tilde{k}$, we employ two measures: the Neumann data loss and the relative error of $k$. The Neumann data loss, defined in (3) with $\omega=\partial\Omega$ (fully observed setting), measures the difference between the predicted and observed Neumann data. Meanwhile, the relative error provides a normalized measure of the difference between $\tilde{k}$ and the exact coefficient $k$:
where $\|\cdot\|_{L^{2}(\overline{\Omega})}$ denotes the $L^{2}$ norm in $\overline{\Omega}$.
Figure 3 presents the Neumann data loss (left) and the relative error of $k$ (right) side by side on a logarithmic scale, where each point represents the averaged value over a 10 epoch interval. This averaging smooths out short-term fluctuations, providing a clearer view of the overall training dynamics.
Figure 3: (Left) Neumann training and validation loss (3) with $\omega=\partial\Omega$, and (Right) relative error of $k$ over epochs (log scale). Each point represents the averaged value over a 10 epoch interval.
As shown in the left plot of Figure 3, the training loss (blue dashed line) decreases sharply at the beginning of the training process and continues to decrease, reaching around $10^{-4}$ during 20–50 epochs. After 50 epoch, when the learning rate is reduced, the training loss continues to decrease and finally stabilizes around $10^{-6}$. The validation loss (red line) shows larger oscillations due to the smaller validation set size and the fact that it is computed once per training epoch, which increases its sensitivity to parameter updates. Despite this variability, the validation loss also experiences a significant decrease after 50 epoch, reaching approximately $10^{-8}$. These observations suggest that the model makes effective use of the Neumann boundary data, with the learning rate schedule playing a crucial role in refining the parameter updates. On the right, the relative error of $k$ follows a similar trend as the Neumann training loss (blue dashed line on the left). After a rapid initial decrease, it remains around $10^{-3}$ during 20-50 epochs, then undergoes a sharp decrease and stabilizes approximately $10^{-4}$. This simultaneous decrease in both metrics indicates that reducing the Neumann data loss effectively guides $\tilde{k}$ to the exact coefficient $k$.
These results support the claim in Problem that $k$ can be uniquely determined from Cauchy data.
@@ 4.1 Example 1: Constant Coefficient
(c) absolute error $|k-\tilde{k}|$
@@
Figure 4: Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$.
Figure 4 visualizes the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$, providing a clear assessment of the reconstruction accuracy. The results show that the proposed method yields a highly accurate approximation of $k$ with only minor deviations. This confirms the effectiveness of the approach in reconstructing constant coefficients with minimal error.
Up to this point, our analysis of the Neumann data loss and reconstruction error has demonstrated that the model can accurately reconstruct a constant coefficient. These numerical results confirm that training based on Neumann data plays a crucial role in optimizing the model. To provide a more intuitive understanding of how the boundary conditions operate in practice, we now examine the operation of the DtN map as illustrated in Figure 5.
[…]
Figure 5: Illustrations of the Dirichlet-to-Neumann (DtN) map $\Lambda$ for a 2D problem. The solution $p(x,y)$ is shown as a heat map. The Dirichlet boundary values $p(x,1)$ and $p(1,y)$ are mapped to their corresponding Neumann data via the DtN map.
4.2 Example 1-2: Constant Coefficient with Partially Observed Data
In this example, we address the inverse problem of the partially observed data. In many real-world scenarios, measurements may be unavailable in certain regions due to sensor limitations or other constraints.
We consider the case of a constant coefficient $k=1$. The computational domain defined as the unit square $\Omega=(0,1)\times(0,1)$, represented by a mesh with 66,049 points. To simulate partial observations, we assume that measurements are unavailable in certain parts of the boundary, covering approximately 20% of $\partial\Omega$. Specifically, we define the region $\Gamma_{0}$, where measurements are unavailable, as follows: measurements are unavailable for $y\in[\frac{3}{5},\frac{4}{5}]$ on the left boundary ($x=0$), for $y\in[\frac{1}{5},\frac{2}{5}]$ on the right boundary ($x=1$), for $x\in[\frac{2}{5},\frac{3}{5}]$ on the bottom boundary ($y=0$), and for $x\in[\frac{3}{5},\frac{4}{5}]$ on the top boundary ($y=1$). In these regions, both the Dirichlet boundary condition and the source function are set to zero; that is, $g=0$ and $f=0$ on $\Gamma_{0}$. The remaining boundary, denoted by $\Gamma=\partial\Omega\setminus\Gamma_{0}$, is where measurements are available. On $\Gamma$, the Dirichlet boundary condition is defined as $g(x,y)=\cos(a\pi x)\cos(b\pi y)$, where $a$ and $b$ are independently sampled from a uniform distribution over $[0,2]$. Similarly, the source function is defined on $\Gamma$ as $f(x,y)=(a^{2}+b^{2})\pi^{2}\cos(a\pi x)\cos(b\pi y)$. The Neumann data is measured only on $\Gamma$, and the Neumann data loss function (3) is defined with $\omega=\Gamma$.
Figure 6: (Left) Boundary regions and (Right) Dirichlet boundary condition $g$.
Figure 7: Relative error of $k$ over epochs (log scale) for both fully and partially observed settings. Each point represents the average over a 10-epoch interval.
(a) exact $k$
(b) reconstructed $\tilde{k}$
(c) absolute error $|k-\tilde{k}|$
Figure 8: Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$ in the partially observed setting.
Figure 7 shows the relative error of $k$ on a logarithmic scale, comparing the cases of fully observed ($\omega=\partial\Omega$) and partially observed ($\omega=\Gamma$). Despite the limited observations in the partially observed setting, the relative error shows a convergence trend similar to that of the fully observed setting. This suggests that the proposed method remains robust even when only partial observations are available, achieving comparable accuracy. Figure 8 presents the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error $|k-\tilde{k}|$, providing a visual assessment of the reconstruction performance in the partially observed setting. The results indicate that the proposed method accurately reconstructs the coefficient even with partial boundary measurements, demonstrating its robustness in handling incomplete data.
4.3 Example 2: Spatially Varying Coefficient
In this example, we test the algorithm with a spatially varying function $k$. In this case, we set
$k(x,y)=0.9\sin\left(\frac{\pi}{3}(x+y+0.1)\right).$
All the other setups including the computational domain, number of mesh points, and Dirichlet boundary condition, are the same as the previous Example 1. The source function is defined as $f(x,y)=\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$ by computing the divergence of the flux associated with the given function.
(a) exact $k$
(b) reconstructed $\tilde{k}$
(c) absolute error $|k-\tilde{k}|$
Figure 9: Illustrations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three dimensions, and the absolute error $|k-\tilde{k}|$ in a heat map.
@@ 4.3 Example 2: Spatially Varying Coefficient
Figure 9 illustrates the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in a three-dimensional space, and their absolute error in a heat map, providing a comparative view of their structures. The reconstructed coefficient $\tilde{k}$ closely matches the exact $k$, effectively capturing its spatial variations, with the absolute error remaining at a low level.
@@
Figure 10: Slices of the coefficient $\tilde{k}$ for the case of $k=0.9\sin\left(\dfrac{\pi}{3}(x+y+0.1)\right)$.
Figure 10 shows the stability of the reconstruction process by comparing both the boundary and interior regions. Over the entire domain, the reconstructed coefficient $\tilde{k}$ closely follows the behavior of $k$, showing strong agreement both at the boundaries and the interior. This consistency highlights the robustness of the method in accurately capturing spatial variations.
@@ 4.4 Example 2-2: Spatially Varying Coefficient with Partially Observed Data
In this example we consider the case of spatially varying coefficients on the partially observed data.
The coefficient $k$ is defined as $k(x,y)=0.9\sin(\frac{\pi}{3}(x+y+0.1))$. All other setups including the computational domain, number of mesh point, Dirichlet condition, and loss function are the same as the previous Example 1-2. The source function is defined as $f(x,y)=\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$ by computing the divergence of the flux associated with the given function. This function is imposed on $\Gamma$, while $f$ is set to zero on $\Gamma_{0}$.
@@
(c) absolute error $|k-\tilde{k}|$
Figure 11: Visualizations of the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three dimensions, and the absolute error $|k-\tilde{k}|$ in a heat map in the partially observed setting.
Figure 11 presents the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$ in three-dimensional space, and the absolute error $|k-\tilde{k}|$ in a heat map, providing a visual comparison of the reconstruction performance under partially observed setting. The results confirm the robustness of the proposed method in accurately reconstructing the coefficient even with limited boundary measurements.
@@
Constant Fully $1.45\times 10^{-12}$ $1.05\times 10^{-6}$ 118.38 0.8346
Partially $1.38\times 10^{-8}$ $1.00\times 10^{-3}$ 78.61 0.9438
Spatially varying Fully $2.65\times 10^{-6}$ $1.50\times 10^{-3}$ 54.85 0.999996
Partially $2.47\times 10^{-7}$ $3.80\times 10^{-4}$ 65.16 0.999997
Table 1: Comparison of MSE, MAE, PSNR, and SSIM for the constant and spatially varying cases under fully and partially observed settings.
Table 1 presents the mean squared errors (MSE) for the results of Examples 1, 1-2, 2, and 2-2, comparing the fully and partially observed settings for both constant and spatially varying cases. In the constant case, the model achieves extremely low reconstruction error, with an MSE of $1.45\times 10^{-12}$ in the fully observed setting. Even in the partially observed setting, the MSE remains low at $1.37\times 10^{-8}$, indicating strong reconstruction performance despite incomplete boundary data. In the spatially varying case, the reconstruction is more challenging due to the spatial variability of the coefficient $k$. Nonetheless, the model achieves a low MSE of $2.65\times 10^{-6}$ in the fully observed setting, demonstrating its capability to approximate the target function accurately. Interestingly, the MSE in the partially observed setting is lower than in the fully observed setting. This may be attributed to the fact that both models converge to highly accurate solutions with very small errors, where minor numerical variations can lead the partially observed setting to yield a slightly lower MSE.

### v3

@@ 3.4 Loss Functions
The loss function based on Problem is defined as
@@
where $\mathcal{D}=\{(g_{i},h_{g_{i}})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g_{i}\subset\omega,h_{g_{i}}=k\partial_{\nu}p_{g_{i}}|_{\omega}\}$, $\tilde{\mathcal{D}}=\{(g_{i},\tilde{k}\partial_{\nu}\tilde{p}_{g_{i}}|_{\omega})_{i=1,\dots,N}\mid g_{i}\in H^{1/2}(\partial\Omega),\,\operatorname{supp}g_{i}\subset\omega\}$, and $\|\cdot\|_{L^{2}(\omega)}$ denotes the $L^{2}$ norm on $\omega$. The loss function $\mathcal{L}$ is minimized by adjusting the model parameters $\boldsymbol{\theta}$. If the loss function reaches zero, then $\mathcal{D}=\tilde{\mathcal{D}}$.
@@
$\frac{\partial\mathcal{L}}{\partial\boldsymbol{\theta}}=\frac{\partial\mathcal{L}}{\partial\tilde{k}_{\text{fem}}}\cdot\frac{\partial\tilde{k}}{\partial\boldsymbol{\theta}}.$ (3)
Here, $\tilde{k}$ represents the neural network output as a PyTorch tensor and $\tilde{k}_{\text{fem}}=\tilde{k}.\texttt{detach()}.\texttt{numpy()}$ is a coefficient that shares the same values as $\tilde{k}$ but is converted to a NumPy array for use in the forward PDE solve within FEniCS. The detach() operation disconnects $\tilde{k}$ from PyTorch’s autograd graph, meaning that any gradient information associated with it is lost beyond this point. As a result, the gradient $\partial\mathcal{L}/\partial\tilde{k}_{\text{fem}}$ is computed efficiently using the adjoint method in FEniCS, while $\partial\tilde{k}/\partial\boldsymbol{\theta}$ is obtained via PyTorch’s automatic differentiation [Paszke19]. Multiplying these two terms yields the desired gradient $\partial\mathcal{L}/\partial\boldsymbol{\theta}$.
4 Numerical Experiments
In this section, we present several numerical experiments to validate and demonstrate the effectiveness of the proposed framework introduced in Section 3. The following setup is common across all examples. The neural network is trained for 100 epochs with a batch size of 1, using PyTorch [Paszke19] and the ADAM optimizer [Kingma14], a variant of stochastic gradient descent. All experiments are performed on a system equipped with an NVIDIA GeForce RTX 3090 GPU (24GB VRAM), an Intel Core i9-12900K CPU, and 62GB of RAM. The software environment includes CUDA 11.8 and NVIDIA driver version 522.06.
4.1 Example 1-1: Constant Coefficient
In this example, we consider a constant coefficient $k$ to provide a baseline for evaluating model performance under simple conditions. Specifically, we set $k=1$. The computational domain is the unit square $\Omega=(0,1)\times(0,1)$, discretized using a triangular mesh with 4,225 degrees of freedom for the linear Lagrange finite element space. The Dirichlet boundary condition is defined as $g(x,y)=\cos(a\pi x)\cos(b\pi y)|_{\partial\Omega}$, where $a$ and $b$ are independently sampled from the uniform distribution over $[0,2]$. This experiment uses a Cauchy dataset containing 2,048 samples. The source function is given by $f(x,y)=-\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$. The learning rate is initialized at 0.001 and reduced by a factor of 0.25 every 20 epochs to ensure stable convergence and prevent overshooting in the later stages of training.
To evaluate the effectiveness of our method in reconstructing the coefficient $k$, we employ two measures: the Neumann data loss and the relative error of $\tilde{k}$. The Neumann data loss, defined in (2) with $\omega=\partial\Omega$ (fully observed setting), measures the difference between the predicted and observed Neumann data. Meanwhile, the relative error is defined as
where $\|\cdot\|_{L^{2}(\overline{\Omega})}$ denotes the $L^{2}$ norm in $\overline{\Omega}$, and provides a normalized measure of the difference between $\tilde{k}$ and the exact coefficient $k$.
(a) Neumann data loss and relative error of $\tilde{k}$
(b) absolute error |kk~||k-$\tilde${k}|.
Figure 3: Example 1-1. Training dynamics and reconstruction accuracy. (a) Neumann data loss (blue, left axis) and relative error of $\tilde{k}$ (red, right axis) over training epochs (log scale). Each point represents the averaged value over a 10-epoch interval. (b) Absolute error distribution between the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$.
Figure 3 illustrates both the training dynamics and the reconstruction accuracy. Figure 3(a) shows the Neumann data loss (blue, left axis) and the relative error of $\tilde{k}$ (red, right axis), both plotted on a logarithmic scale. Each point corresponds to the average over a 10-epoch interval, which smooths out short-term fluctuations and highlights the overall decreasing trend. The simultaneous reduction of both metrics demonstrates that minimizing the Neumann data loss effectively drives $\tilde{k}$ toward the exact coefficient $k$, supporting the uniqueness claim in Problem. Figure 3(b) visualizes the absolute error |kk~||k-$\tilde${k}| between the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$. The results confirm that the proposed method provides a highly accurate approximation of $k$, with only minor deviations.
To provide a more intuitive understanding of how the boundary conditions operate in practice, we now examine the operation of the DtN map as illustrated in Figure 4.
@@ 4.1 Example 1-1: Constant Coefficient
Figure 4: Example 1-1. Illustration of the Dirichlet-to-Neumann (DtN) map. The solution $p(x,y)$ is shown and the Dirichlet boundary values $p(x,1)$ and $p(1,y)$ mapped to their corresponding Neumann data via the DtN map are presented.
4.2 Example 1-2: Constant Coefficient with Partially Observed Data
In this example, we address the inverse problem under the partially observed data. In many real-world scenarios, measurements may be unavailable in certain regions due to sensor limitations or other constraints.
We consider the case of a constant coefficient $k=1$. All other configurations, including the computational domain, degrees of freedom, the number of data samples, and the learning rate, are the same as in Example 1-1. To simulate partial observations, we assume that measurements are unavailable in certain parts of the boundary, covering approximately 40% of $\partial\Omega$. Specifically, we define the region $\Gamma_{0}$, where measurements are unavailable, as follows: measurements are unavailable for $y\in[\frac{2}{5},\frac{4}{5}]$ on the left boundary ($x=0$), for $y\in[\frac{1}{5},\frac{3}{5}]$ on the right boundary ($x=1$), for $x\in[\frac{1.5}{5},\frac{3.5}{5}]$ on the bottom boundary ($y=0$), and for $x\in[\frac{2.5}{5},\frac{4.5}{5}]$ on the top boundary ($y=1$). In these regions, both the Dirichlet boundary condition and the source function are set to zero; that is, $g=0$ and $f=0$ on $\Gamma_{0}$. The remaining boundary, denoted by $\Gamma=\partial\Omega\setminus\Gamma_{0}$, is where measurements are available. On $\Gamma$, the Dirichlet boundary condition is defined as $g(x,y)=\cos(a\pi x)\cos(b\pi y)$, where $a$ and $b$ are independently sampled from a uniform distribution over $[0,2]$. Similarly, the source function is defined on $\Gamma$ as $f(x,y)=-\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$. The Neumann data is measured only on $\Gamma$, and the Neumann data loss function (2) is defined with $\omega=\Gamma$.
Figure 5: Example 1-2. (Left) Boundary regions and (Right) Dirichlet boundary condition $g$. The values for which $g=0$ indicate that no data is available.
(a) Relative error of $\tilde{k}$ over epochs (log scale) for fully and partially observed settings.
(b) absolute error |kk~||k-$\tilde${k}|.
Figure 6: Example 1-2. Training results and reconstruction accuracy for both fully and partially observed settings. (a) Relative error curves of $\tilde{k}$ under different observation settings. (b) Absolute error |kk~||k-$\tilde${k}| in the partially observed case.
Figure 6 illustrates both the training results and the reconstruction accuracy. Figure 6(a) shows the relative error of $\tilde{k}$ on a logarithmic scale, comparing the fully observed ($\omega=\partial\Omega$) and partially observed ($\omega=\Gamma$) settings. Although the relative error in the partially observed setting is slightly higher than that in the fully observed case, it steadily decreases to a low value, indicating effective reconstruction. Figure 6(b) visualizes the absolute error |kk~||k-$\tilde${k}| between the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ in the partially observed setting. These results indicate that the proposed method can accurately reconstruct the coefficient even with partial boundary measurements, demonstrating its robustness in handling incomplete data.
4.3 Example 2-1: Spatially Varying Coefficient
In this example, we test the algorithm with a spatially varying function $k$. Specifically, we set
$k(x,y)=0.9\sin\left(\frac{\pi}{3}(x+y+0.1)\right).$
All other configurations, including the computational domain, degrees of freedom, Dirichlet boundary condition, number of data samples, and learning rate, are the same as in Example 1-1. The source function is defined as $f(x,y)=-\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$ by computing the divergence of the flux associated with the given function.
@@ 4.3 Example 2-1: Spatially Varying Coefficient
(c) absolute error |kk~||k-$\tilde${k}|
@@
Figure 7: Example 2-1. Visualizations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ using 3D surface plots, along with the absolute error |kk~||k-$\tilde${k}| shown as a 2D heat map.
@@
Figure 7 illustrates the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and their absolute error, providing a comparative view of spatial structures. The reconstructed coefficient $\tilde{k}$ closely matches the exact $k$, effectively capturing its spatial variations, with the absolute error remaining at a low level.
@@
Figure 8: Example 2-1. Slices of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ along selected lines in the domain.
Figure 8 shows the reconstruction results by comparing both the boundary and interior regions. Over the entire domain, the reconstructed coefficient $\tilde{k}$ closely follows the behavior of $k$, showing strong agreement both at the boundaries and the interior. This consistency highlights the robustness of the method in accurately capturing spatial variations.
@@ 4.4 Example 2-2: Spatially Varying Coefficient with Partially Observed Data
In this example, we consider the case with partially observed data. The coefficient $k$ is defined as $k(x,y)=0.9\sin(\frac{\pi}{3}(x+y+0.1))$. All other configurations, including the computational domain, degrees of freedom, Dirichlet boundary condition, number of data samples, and learning rate, are the same as in Example 1-2. The source function is defined as $f(x,y)=-\nabla\cdot(k(x,y)\nabla(\cos(a\pi x)\cos(b\pi y)))$ by computing the divergence of the flux associated with the given function. This function is imposed on $\Gamma$, while $f$ is set to zero on $\Gamma_{0}$.
Figure 9: Example 2-2. Relative error of $\tilde{k}$ over epochs (log scale) for both fully and partially observed settings. Each point represents the average over a 10-epoch interval.
Figure 9 shows the relative error of $\tilde{k}$ on a logarithmic scale, comparing the fully observed and partially observed settings. As in the constant coefficient case, the relative error under the partially observed setting is slightly higher, but it still decreases steadily, indicating good convergence behavior.
@@
(c) absolute error |kk~||k-$\tilde${k}|
Figure 10: Example 2-2. Visualizations of the exact coefficient $k$ and the reconstructed coefficient $\tilde{k}$ using 3D surface plots, along with the absolute error |kk~||k-$\tilde${k}| shown as a 2D heat map in the partially observed setting.
Figure 10 presents the exact coefficient $k$, the reconstructed coefficient $\tilde{k}$, and the absolute error |kk~||k-$\tilde${k}|, providing a visual comparison of the reconstruction performance under the partially observed setting. The results confirm the robustness of the proposed method in accurately reconstructing the spatially varying coefficient, even with limited boundary measurements.
@@
Constant Fully $3.10\times 10^{-11}$ $4.11\times 10^{-6}$ 105.09 0.999999999
Partially $8.93\times 10^{-10}$ $2.11\times 10^{-5}$ 90.49 0.999999997
Spatially varying Fully $95\times 10^{-8} 1.1$$11\times 10^{-4}$ 76.18 0.999999521
Partially $4.10\times 10^{-8}$ $1.41\times 10^{-4}$ 72.95 0.999999575
Table 1: Example 1-1 - Example 2-2. Comparison of MSE, MAE, PSNR, and SSIM for the constant and spatially varying cases under fully and partially observed settings.
Table 1 presents the performance metrics—mean squared error (MSE), mean absolute error (MAE), peak signal-to-noise ratio (PSNR), and structural similarity index (SSIM)—for Examples 1-1, 1-2, 2-1, and 2-2. It compares the reconstruction quality under fully and partially observed settings for both the constant and spatially varying coefficient cases. PSNR measures the numerical similarity of the reconstructed values to the exact values [Huynh08], while SSIM evaluates the structural consistency between the reconstructed and exact values [Wang04]. Higher values of PSNR and SSIM indicate better reconstruction performance, with SSIM values closer to 1 reflecting greater structural similarity. In contrast, lower values of MSE and MAE correspond to smaller reconstruction errors.
In the constant case, the fully observed setting outperforms the partially observed setting in all metrics, including MSE, MAE, PSNR, and SSIM. Nevertheless, the partially observed setting achieves a very low MSE of $8.93\times 10^{-10}$ and a SSIM value close to 1, indicating that the overall reconstruction remains highly accurate. In the spatially varying case, all metrics are slightly lower than in the constant case, reflecting the increased complexity of the coefficient. The fully observed setting yields lower MSE and MAE values and achieves a higher PSNR, indicating a more accurate reconstruction. Although the SSIM value is slightly higher in the partially observed setting, the difference is negligible.